The perimeter, P, of a rectangular concrete slab is 46 m and it's area, A, is 90 m^2. Use the formula P + 2l + 2w. Determine the dimensions of the slab. Show your work.
I have tried to do this and keep getting the most ridiculous answers. This question has really stumped me and I have come to a stand still.

Answer by jim_thompson5910(35256)

A = LW ... area of rectangle

90 = LW

LW = 90

L = 90/W

------------------------------------

P = 2L + 2W

46 = 2L + 2W

46 = 2(90/W) + 2W

46 = 180/W + 2W

W*(46) = W(180/W + 2W) ... Multiply both sides by W so we can clear out that fraction.

W*(46) = W(180/W) + W(2W)

46W = 180 + 2W^2

0 = 180 + 2W^2 - 46W

2W^2 - 46W + 180 = 0

2(W^2 - 23W + 90) = 0

W^2 - 23W + 90 = 0/2

W^2 - 23W + 90 = 0

(W - 18)(W - 5) = 0

W - 18 = 0 or W - 5 = 0

W = 18 or W = 5

if the width is W = 18, then the length is L = 90/W = 90/18 = 5.

if the width is W = 5, then the length is L = 90/W = 90/5 = 18.


So regardless of which value you pick for W, the other value for L will be the other solution.

In essence, the dimensions are 18 meters by 5 meters


Check:

A = LW = 18*5 = 90 square meters ... area checks out

P = 2L + 2W = 2*18 + 2*5 = 36+10 = 46 meters ... perimeter checks out